OpenStudy (highschoolmom2010):

jasmine made a cylindrical vase in which the sum of the lateral area and area of one of the bases is 3000 square centimeters. the vase was 50 centimeters. find the radius. explain the method you used.

ganeshie8 (ganeshie8):

\(2\pi r h + \pi r^2 = 3000\) \(h = 50\)

ganeshie8 (ganeshie8):

substitute vaue of \(h\) in first equation

ganeshie8 (ganeshie8):

\(2\pi r*50 + \pi r^2 = 3000\) \(100\pi r + \pi r^2 = 3000\)

ganeshie8 (ganeshie8):

solve it using quadratic formula

ganeshie8 (ganeshie8):

wolfram says \(r \approx 8.78\)

ganeshie8 (ganeshie8):

let me knw if smthn doesnt make sense

OpenStudy (highschoolmom2010):

thats what i got but my teacher says it is incomplete :/

OpenStudy (phi):

yes, but you should fix the fourth line in your write-up. You wrote 12000 pi^2 inside the square root sign and it should say 12000 pi (I assume it is a "typo", because you got the correct number on the next line) also, on that same line, you should extend the line to show that the 2pi also divides the -100 pi